OpenStudy (anonymous):

PLEASE HELP? Compare and contrast, in complete sentences, the difference between the products of y2y5 and (y2)5. Provide the expanded form and product of each.

OpenStudy (anonymous):

do you mean these two: \(\large y^2y^5 \) and \(\large (y^2)^5 \) ???

OpenStudy (anonymous):

Yes :) @dpaInc

OpenStudy (anonymous):

ok.. what does \(\large y^2 =\) ???

OpenStudy (anonymous):

similarly, \(\large y^5 = \) ????

OpenStudy (anonymous):

I don't know... @dpaInc

OpenStudy (anonymous):

what does \(\large y\cdot y= \) ???

OpenStudy (anonymous):

y^2?

OpenStudy (anonymous):

ok... ill give you the first one... \(\large y^2 = y\cdot y \) so \(\large y^5 = y\cdot y\cdot y\cdot y\cdot y \) still with me? your answer is correct in the post just above this one....

OpenStudy (anonymous):

Yeah i'm with you

OpenStudy (anonymous):

so, \(\large y^2y^5 = \color {red}{y^2}\cdot \color {blue}{y^5} = \color {red}{y\cdot y}\cdot \color {blue}{y\cdot y\cdot y\cdot y\cdot y} = y^7 \)

OpenStudy (anonymous):

now let's do \(\large (y^2)^5 \)

OpenStudy (anonymous):

So that is the extended version of the first one?

OpenStudy (anonymous):

yes...^^^

OpenStudy (anonymous):

okay i got it

OpenStudy (anonymous):

so for the second one..

OpenStudy (anonymous):

\(\large (y^2)^5 =(y^2)\cdot (y^2)\cdot (y^2)\cdot (y^2)\cdot (y^2)\) can you expand what's in each of those parenthesis?

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

(y^2)^5 = (y times y) (y times y) (y times y) (y times y) (y times y)

OpenStudy (anonymous):

yes.... so using exponents, what does (y^2)^5 = ???

OpenStudy (anonymous):

So for both of them i got (y^2)^5 = y^2 * y^2 * y^2 * y^2 * y^2 y^2y^5 = y^2 * y^5 = y * y * y * y * y = y^7

OpenStudy (anonymous):

what would the product of the 2nd one be?

OpenStudy (anonymous):

(y^2)^5 = y^2 * y^2 * y^2 * y^2 * y^2 = y * y * y * y * y * y * y * y * y * y = y^10

OpenStudy (anonymous):

I have to compare and contrast them as well

OpenStudy (anonymous):

the purpose ( i think) of this problem is to show you a very important property of exponents... \(\large y^2y^5=y^{2+5}=y^7\) \(\large (y^2)^5=y^{2\cdot 5}=y^{10}\)

OpenStudy (anonymous):

Okay i think i got it all, thank you so much @dpaInc ill have more questions in a few :)

OpenStudy (anonymous):

yw...